OpenStudy (anonymous):

let F(x)=-4x+5 and G(x)=-3x-2 find (f o g)(3)

OpenStudy (precal):

f(g(3)) g(3)=-3(3)-3=-9-3=-12 g(3)=-12 f(-12)=-4(-12)+5=48+5=53

OpenStudy (anonymous):

thank you